OpenStudy (alexisreynosaboi):

A standard deck of 52 playing cards contains 13 cards in each of four suits: diamonds, hearts, clubs, and spades. Two cards are chosen from the deck at random. What is the approximate probability of choosing one club and one heart? 0.0588 0.0637 0.1176 0.1275

OpenStudy (kevin):

1 deck has 13 clubs and 13 hearts So total clubs/total cards x total hearts/total cards

OpenStudy (dallascowboys88):

b

OpenStudy (jango_in_dtown):

Let A denote the event of choosing one club. then P(A)=13/52. and B denote the event of choosing one heart when one club is already chosen. Then P(B)=13/51. Then required probability =2C1*P(A .B)=2*P(A)P(B) =2*(13/52)(13/51)=26/51=0.1275 approx

OpenStudy (jango_in_dtown):

2C1 because any of them can be chosen first...